OpenStudy (anonymous):

On the first day of ticket sales the school sold 2 senior citizen tickets and 1 student ticket for a total of 21$. The school took in 129$ on the second day by selling 11 senior citizen tickets and 10 student tickets. Find the price of a senior citizen ticket and a student ticket.

OpenStudy (campbell_st):

ok... so you have 2 equations let C = senior citizens S = students so ticket sales 2C + S = 21 (day 1) 11C + 10S = 129 (day 2) take the day 1 equation and write it with S as the subject S = 21 - 2C now substitute this into the day 2 says information 11C + 10(21 - 2C) = 129 after distrubuting you get 11C + 210 - 20C = 129 subtract 210 from both sides and collect the like terms -9C = -81 I'll let you solve to equation to find the cost of a senior citizen ticket. Hope this helps

OpenStudy (campbell_st):

to find the price of a student ticket... substitute the value of C into S = 21 - 2C

OpenStudy (anonymous):

I'm lost. how would you solve the first equation without a common variable..

OpenStudy (anonymous):

I got 9 for senior and 2 student

OpenStudy (campbell_st):

I think the student is 3 and the senior citizen is 9

OpenStudy (anonymous):

Ah yes i added wrong